OpenStudy (anonymous):

A certain keyboard has five more than twice as many keys across as it does vertically. This special keyboard has a total of 168 keys. Find the amount of keys positioned vertically and across.

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@courtney1967 show me how to figure

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

Lets draw a picture. Can you show me, first, what 5 more than twice a number looks like in numbers. Say we call x the number.

OpenStudy (anonymous):

5+2x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

Now we know that the total number of keys are 168 right?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

So 5 more than twice a number of keys gives us 168 keys. What equation can you write that shows this?

OpenStudy (anonymous):

5+2x=168

OpenStudy (anonymous):

Dude.. rock on!

OpenStudy (anonymous):

Alright, let's see what x is.

OpenStudy (anonymous):

yess feels good to know it

OpenStudy (anonymous):

How would you find out what x is in 5+2x=168?

OpenStudy (anonymous):

-5 from 168 and 5

OpenStudy (anonymous):

or divide?

OpenStudy (anonymous):

by 2

OpenStudy (anonymous):

Right! So we subtract 5 from both sides. 2x = 168 - 5 2x = 163 So now what?

OpenStudy (anonymous):

divide by 2

OpenStudy (anonymous):

but 2 doesent go in 163

OpenStudy (anonymous):

81.5

OpenStudy (anonymous):

Your right, so we made a mistake. Do you see where?

OpenStudy (anonymous):

by subtracting 168-5

OpenStudy (anonymous):

We said that 2x + 5 was equal to 168 keys, but we needed to look at it this way. 

OpenStudy (anonymous):

it 80

OpenStudy (anonymous):

its

OpenStudy (anonymous):

thats the height

OpenStudy (anonymous):

I mean

OpenStudy (anonymous):